# Finding Unexpected Non-Helical Tracks

### v1

@@ I Introduction
Particle colliders such as the Large Hadron Collider (LHC) have tremendous potential to discover new physics ATLAS:2012yve; CMS:2012qbp; annurev:/content/journals/10.1146/annurev-nucl-101920-014923; annurev:/content/journals/10.1146/annurev-nucl-102419-052854; LHCb:2015yax; LHCb:2014vgu; ATLAS:2010isq; CMS:2010ifv, but require computationally expensive analysis to take full advantage of their data. One of the most challenging tasks is to identify a set of detector responses due to an individual particle, a track, among the enormous number left by all particles produced in a collision. Considering every possible subset is computationally intractable task due to the vast combinatorics in the most general case Strandlie:2010zz. Traditionally, simplifying assumptions are employed to make the task feasible ATLAS:2023iat. Primary among these is restricting the search to tracks which follow the helical trajectory expected of electrically-charged Standard Model particles in a uniform magnetic field. This greatly simplifies the track identification task, but leaves experiments incapable of finding non-helical tracks, shrinking the potential for discovery.
@@
Several theories of physics beyond the Standard Model (BSM) have interactions which produce non-helical trajectories, such as magnetic monopoles ATLAS:2019wkg; CDF:2005cvf or quirks Kang:2008ea; Evans:2018jmd; Knapen_2017. Recent work on quirks  Sha:2024hzq has demonstrated the capability to efficiently find non-helical tracks when the predicted trajectories are specified in advance. However, the space of possible particle trajectories is much larger than what is described by the current set of BSM theories. Model-agnostic tracking, capable of finding tracks without a pre-existing theory or pre-specified trajectory, would broaden experimental capacity for unexpected discoveries, especially in low-background, high-fidelity environments such as tracking detectors PhysRev.43.491. Tracks with striking, surprising trajectories may exist in current experimental datasets, undiscovered only because they are invisible to current algorithms.
@@
Most traditional track-finding algorithms ATLAS:2017kyn perform alternating track parameter fitting and iterative hit finding, which requires knowledge of the explicit parametric form of the track, and so are fundamentally unsuited to model-agnostic tracking. However, recent developments in machine learning such as graph neural networks (GNNs) Bronstein:2016thv allow for track-finding algorithms where the form of the track is never specified explicitly, only implicitly via the sample used to train the algorithm ExaTrkX:2020nyf; Ju_2021. While not the original goal of ML tracking, this new approach separates the finding and fitting steps Caillou:2024smf; Caillou:2022hly, and therefore allows for finding non-helical trajectories when trained on such examples Sha:2024hzq. Crucially, it may also open the door to model-agnostic tracking, if the broad space of possible trajectories can be described implicitly via a training sample, and the tracker learns to generalize beyond the specific training examples.
@@
In Section II, the simulation of tracks is described, including the detector geometry and model-agnostic track generation. In Section III, the tracking pipeline is described. Section IV details the reconstruction performance on various sets of tracks, including background rejection and tests of generalization. Section V discusses the results and next steps.
@@ II.1 Detector Geometry
Particle energy loss and hit location uncertainty are neglected, but similar work did not show strong sensitivity Sha:2024hzq.
@@ II.2 SM Background
Madgraph5 Alwall:2014hca is used to generate $pp\rightarrow t\bar{t}$ events, which are showered and hadronized using Pythia8 Bierlich:2022pfr. Initial velocity and momenta are used to propagate Standard Model particles in helical trajectories under the assumption that there is no energy loss or secondary scattering.
All events use a pileup of $\mu=1$. The impact of pile-up on performance is not yet known, but in other studies the same GNN pipeline performed well in high pile-up conditions similar to those expected at the high-luminosity LHC ExaTrkX:2021abe. Such performance may extend to the case of generalized smooth tracks, or it may be more susceptible to confusion in such hit dense environments. A rigorous study is out of scope for this initial proof-of-principle.
Refer to captionRefer to caption
@@ II.3 Non-Helical Track Generation
The model-agnostic search aims to be sensitive to a very broad range of possible new particles with unexpected trajectories through the detector. However, it cannot accept every possible trajectory, or all sets of hits would be candidates. Instead, particle tracks are required to follow smooth paths, those which can be parameterized by a curve, $\vec{x}(t)$ where all orders of spatial derivatives with respect to time, $\frac{d^{n}x}{dt^{n}}$, exist and are finite, which accepts all particles with physical paths, where the net force and its derivatives, $m\frac{d^{n}\ddot{x}}{dt^{n}}$, are physical [^1] While this may be untrue for idealized or simplified systems which exhibit discontinuities in external forces, fully physical classical systems disregarding quantum effects such as decays, are subject to classical field interactions, which are infinitely differentiable with respect to position and time Landau:1975pou.. It rejects candidates whose paths are discontinuous or kinked. This method of track generation narrows our analysis by disregarding non-prompt decays, and particles that scatter off of detector material. While many particles exhibit non-smooth behavior in detector environments, a smoothness assumption provides a theoretically sound foundation for demonstrating this proof-of-concept analysis.
Without assumption about the underlying physical theory, three-component spatial curves can be fully described in a Fourier basis by three sequences of Fourier amplitudes, $(a_{n},b_{n},c_{n})$. Such paths, however, are not necessarily smooth unless each sequence approaches zero as $n\rightarrow\infty$ at least as fast as a Schwartz function f(n)f(n)Stein:2003; Hormander:1990; Reed:1980. Explicitly, smoothness is guaranteed if $a_{n},b_{n},c_{n}\leq f(n)$ for all $n$.
@@
For demonstration, non-helical tracks with a small number of frequencies are also prepared, by sampling from finitely many frequencies in Fourier space, where each mode has a maximum amplitude of 60cm. An $N$-frequency track is defined as the collection of points where a function $\vec{g}_{N}({t})=\sum_{i=1}^{3}\sum_{n=-N}^{N}c_{ni}\exp(\frac{2\pi int}{\lambda})\hat{x_{i}}$ intersects the detector layers, where the $c_{ni}$ are all randomly generated with a uniform distribution in the interval $[0,60]$ cm, and $\lambda$ is set equal to the radius of the detector which is a choice further discussed in App. .3.
@@ III Tracking
Many tracking algorithms, such as a Kalman Filter, assume a parametric track path and ionization and interaction properties typical of known particles. An initial set of hits allows for initial estimates of the track parameters, which in turn allows for propagation of the track to subsequent detector surfaces and a calculation of a window within which additional hits should be considered. This hit-by-hit approach naturally weaves together track finding and fitting. Machine learning can be used in a variety of ways to perform track findingduarte2022graph; dezoort2023graph; thais2022graph. For example, ML can be used to propagate initial tracks in a Kalman Filter heinrich2024combined, or more broadly to select the next hit vaage2022reinforcement; kortus2023towards.
The machine-learning-based approach used below does not require an explicit model of the trajectory. Instead, the relationships between hits on the same track are implicitly defined through the example tracks in the training sample. A track is represented as a directed graph, with hits as nodes. Track-finding pipelines, such as Exa.TrkX, which use graph neural networks achieve high track-finding efficiencies in a broad variety of particle experiments, even under high pile-up conditions caillou2022atlas; verma2020particle; hewes2021graph; drielsma2021scalable; akram2022track; jia2024besiii; biscarat2021towards; zdybal2024machine; andrews2021accelerating; dezoort2021charged; Correia_2024; Elabd_2022. The Exa.TrkX pipeline first performs graph construction, where hits are linked into a graph, followed by edge classification, where graph edges are evaluated as their likelihood to connect two hits which belong to the same track, followed by graph segmentation, where edges are pruned and collected hit connections form track candidates. The Exa.TrkX pipeline employed here is modified to relax the requirement that hits are ordered in time according to their distance from the vertex, following Ref. Sha:2024hzq. Graph edges are evaluated by a metric learning model, which learns a mapping from the physical space of hits to an abstract latent space, such that hits on the same track are near each other and hits from other tracks are more distant.
@@
IV Reconstruction Performance
IV.1 Baseline tests
@@
In practice, non-helical tracks are expected to be found among many SM tracks. In the next stage, SM tracks are added to the training sample and given the same labels [^2] In previous work which attempted to learn one specific non-helical track pattern Sha:2024hzq, SM tracks were labeled as fake to suppress their reconstruction. In contrast, here the aim is to learn all possible smooth tracks. Efficiency is lower in the presence of many background tracks, as expected, but remains strong; see Tab. 1. Even in the presence of SM backgrounds, the tracking pipeline can efficiently find non-helical tracks.
IV.2 Generalization tests
The training sample is necessarily finite, and is constructed using a finite number of Fourier modes from one of many Schwartz functions. An important question is whether the tracker has learned the general properties of smooth tracks, allowing it to generalize outside of its training sample. Previous work Sha:2024hzq showed hints that the tracking pipeline may be capable of generalization, learning to reconstruct tracks which fall outside of the space described by the training set.
@@ IV.2 Generalization tests
In one case, training on Schwartz space 5 and testing on 6-5, gives poor performance. The testing set has very long tracks, as shown in Fig. 5, which makes reconstruction under the double matching criterion very difficult. Generalization is possible, but not universal or guaranteed.
IV.3 Dependence on Training Functions
@@
IV.4 Results
The efficiency of the pipeline when trained using Schwartz function 19 is shown in Tab. 3 for a selection of testing samples. The training sample size has been increased to 65,000 events, each with one non-helical track and SM tracks from one SM event. This pipeline has a strong capacity to reconstruct many varieties of smooth tracks when correctly trained, while maintaining a low fake rate.
@@ IV.4 Results
IV.5 Rejection of Standard Model tracks
@@
Helical tracks can be rejected by fitting their points to a helical trajectory and evaluating the goodness of fit, via the $\chi^{2}$. For each track, an initial rapid estimation of the helical parameters is made HANSROUL1988498, followed by an exploration of parameters values in the vicinity, and refinement with the Nelder-Mead method nelder implemented in scipy.minimize 2020SciPy-NMeth. The distribution of the $\chi^{2}$ for SM and non-helical tracks is shown in Fig 7. While some SM tracks will have larger $\chi^{2}$, and some non-helical tracks have low $\chi^{2}$, there is a significant population of non-helical tracks with large $\chi^{2}$ and which are striking by eye. Several examples of fitting helices to hits left by non-helical tracks are shown in Fig. 8.
@@ VI Conclusions
.1 Schwartz Functions
Table 4: Selected Schwartz functions.
Number Expression
1 $f_{1}(n)=60\cdot\exp(-0.01n^{2})$
2 $f_{2}(n)=60\cdot\exp(-0.01(n-7)^{2})$
3 $f_{3}(n)=30\cdot\exp(-0.01n^{2})$
4 $f_{4}(n)=30\cdot\exp(-0.01(n-7)^{2})$
5 $f_{5}(n)=10\cdot\exp(-0.01n^{2})$
6 $f_{6}(n)=10\cdot\exp(-0.01(n-7)^{2})$
7 $f_{7}(n)=45\cdot\exp(-0.01n^{2})$
8 $f_{8}(n)=45\cdot\exp(-0.01(n-7)^{2})$
9 $f_{9}(n)=45\cdot\exp(-0.01(n-3.5)^{2})$
10 $f_{10}(n)=60\cdot\exp(-0.03n^{2})$
11 $f_{11}(n)=45\cdot\exp(-0.02(n-3.5)^{2})$
12 $f_{12}(n)=60\cdot\exp(-0.01(n-3.5)^{2})$
13 $f_{13}(n)=60\cdot\exp(-0.02n^{2})$
14 $f_{14}(n)=52.5\cdot\exp(-0.01n^{2})$
15 $f_{15}(n)=37.5\cdot\exp(-0.01n^{2})$
16 $f_{16}(n)=60\cdot\exp(-0.01(n-1.75)^{2})$
17 $f_{17}(n)=60\cdot\exp(-0.01(n-5.25)^{2})$
18 $f_{18}(n)=60\cdot\exp(-0.015n^{2})$
19 $f_{19}(n)=60\cdot\exp(-0.025n^{2})$
20 $f_{20}(n)=60\cdot\exp(-0.035n^{2})$
21 $f_{21}(n)=60\cdot\exp(-0.025(n-3.5)^{2})$
22 $f_{22}(n)=45\cdot\exp(-0.025(n-7)^{2})$
23 $f_{23}(n)=30\cdot\exp(-0.03(n-7)^{2})$
.2 Appendix: Schwartz Space Properties
It is shown here that the method of smooth track generation detailed in section II.3 is capable of generating every smooth track.
Curves are generated by choosing a Schwartz function, $s\in\mathcal{S}$, as an upper bound for 3 separate sequences of Fourier coefficients, where $\mathcal{S}$ is the set of Schwartz functions. A sequence is chosen for each spatial coordinate, so tracks can be generated that can be parametrized $(f(t),g(t),h(t))$ for $t\in[0,2\pi]$, with periodic components with a period of $2\pi$ and $f(0)=g(0)=h(0)=0$. For each mode of oscillation, $n$, we randomly choose a $c_{n_{x}}$, $c_{n_{y}}$, and $c_{n_{z}}$ in $[0,s(n)]$. To maintain rotational invariance in our Fourier space, we reject all three and choose again if $c^{2}_{n_{x}}+c^{2}_{n_{y}}+c^{2}_{n_{z}}>s^{2}(n)$, so we can think of the three coefficients as being a single point chosen with a uniform distribution inside of some sphere of radius $s(n)$. Thus, we can think of a single track as generated by a sequence of points, where the nth point is chosen to lie inside a sphere of radius $s(n)$. If $f(t)$, $g(t)$, and $h(t)$ are differentiable with respect to time to all orders the track is called ’smooth’.
To show that any smooth track can be generated by choosing the correct Schwartz function, suppose that $g(t)\in\mathbb{R}$ is smooth. We want to show that there exists a function $s\in\mathcal{S}$ for which the $C_{n}\leq s(n)$ for every $n$, where the sequence $C_{n}$ is the Fourier representation of $g$. This can also be stated as:
$C_{n}=\frac{1}{2\pi}\int^{2\pi}_{0}g(x)e^{inx}dx$
We can then integrate by parts to show that
$C_{n}=-\frac{1}{2\pi}\frac{1}{ni}\int^{2\pi}_{0}g^{\prime}(x)e^{inx}dx$
and if we now integrate by parts $k$ times,
$C_{n}=(-1)^{k}\frac{1}{2\pi}(\frac{1}{ni})^{k}\int^{2\pi}_{0}g^{k}(x)e^{inx}dx$
Therefore, it must be the case that
$\left\lvert\int^{2\pi}_{0}g^{k}(x)e^{inx}dx\right\rvert\leq 2\pi\>\text{max}(g^{k}(x))$
for $x\in[0,2\pi]$, and thus it can be immediately shown that
$\left\lvert C_{n}\right\rvert=\left\lvert\frac{1}{2\pi}(\frac{1}{ni})^{k}\int^{2\pi}_{0}g^{k}(x)e^{inx}dx\right\rvert\leq(\frac{1}{n})^{k}\>\text{max}(g^{k}(x))$
Therefore $C_{n}$ is bounded from above by a Schwartz function, and every smooth track can be generated by careful choice of Schwartz function.
It is also important that this method only generates smooth tracks regardless of choice of Schwartz function. This is easy to show, as the Fourier transform is an isomorphism onto the Schwartz space. This means that choosing a sequence of Fourier coefficients to lie on the Schwartz function: $C_{n}=s(n)$, makes the resulting function Schwartz. If instead: $C_{n}\leq s(n)$, then the resulting function is even more rapidly decaying, and is thus still Schwartz, and thus smooth.
It is illustrative to describe helical tracks in a Fourier basis and show that they are a subset of tracks that can be generated by the Schwartz function method described. In tracking, helices are typically fully described by the five parameters $(\phi_{0},p_{\textrm{T}},d_{0},d_{z},\tan(\lambda))$ where $\phi_{0}$ is the initial angle of the particle’s trajectory with respect to the z axis, $p_{\textrm{T}}$ is the transverse momentum, $d_{0}$ is the distance of closest approach to the origin in the xyx-y plane, $d_{z}$ is the distance of closest approach to the origin along the $z$ direction, and $\tan(\lambda)$ is the slope of the helix in the rzr-z plane. The position of a particle for any time $t$ can then be described
[…]
Where $\alpha=\frac{1}{qB}$ is the magnetic field constant.
This form can be expanded in a Fourier basis by first recognizing the form of the x and y components to already be trivially in a Fourier basis and expanding the $z$ component,
$z=d_{z}+\alpha\;p_{\textrm{T}}\tan(\lambda)\;\omega t-\sum_{n=1}^{\infty}\frac{2\;\alpha\;p_{\textrm{T}}\tan(\lambda)}{n}\;\sin{(n\omega t)}$
For any value of $t$ in a compact interval and any values of track parameters, there exists a Schwartz function that generates all the correct coefficients for this helical form. With the assumption that tracks start at the origin, the Schwartz functions chosen in this work will generate this form of helical track if $\alpha\;p_{\textrm{T}}\tan(\lambda))$ is not too large.
.3 Signal track generation
Our signal track generation consists of producing a spatial curve via Fourier series, making a very large number of points on the curve which are evenly spaced by arc length, and recording the point closest to each layer that the curve passes through and labeling it a ”hit”.
When producing the Fourier series, the fundamental wavelength is arbitrarily chosen to be equal in length to the detector radius. So long as sufficiently many frequencies are included in track generation, this will not limit the tracks that can be generated. However, this does put constraints on the tracks, such as those in section IV.1, that are generated with a finite number of frequencies. When making this arbitrary choice it is only important that the fundamental wavelength is a similar length to the radius of the detector, so that the finite Frequency Tracks do not leave the detector too soon or too late.

### v2

@@ I Introduction
Particle colliders such as the Large Hadron Collider (LHC) have tremendous potential to discover new physics Aad et al. (2012); Chatrchyan et al. (2012); Gehrmann and Malaescu (2022); Ferreira da Silva (2023); Aaij et al. (2015, 2014); Aad et al. (2010); Khachatryan et al. (2010), but require computationally expensive analysis to take full advantage of their data. One of the most challenging tasks is to identify a set of detector responses due to an individual charged particle, a track, among the enormous number left by all particles produced in a collision. Considering every possible subset is computationally intractable task due to the vast combinatorics in the most general case Strandlie and Fruhwirth (2010). Traditionally, simplifying assumptions are employed to make the task feasible Aad et al. (2024). Primary among these is restricting the search to tracks which follow the helical trajectory expected of electrically-charged Standard Model particles in a uniform magnetic field. This greatly simplifies the track identification task, but leaves experiments incapable of finding non-helical tracks, shrinking the potential for discovery.
@@
Several theories of physics beyond the Standard Model (BSM) have interactions which produce non-helical trajectories, such as magnetic monopoles Collaboration (2020); Abulencia et al. (2006) or quirks Kang and Luty (2009); Evans and Luty (2019); Knapen et al. (2017). Recent work on quirks  Sha et al. (2024) has demonstrated the capability to efficiently find non-helical tracks when the predicted trajectories are specified in advance. However, the space of possible particle trajectories is much larger than what is described by the current set of BSM theories. Model-agnostic tracking, capable of finding tracks without a pre-existing theory or pre-specified trajectory, would broaden experimental capacity for unexpected discoveries, especially in low-background, high-fidelity environments such as tracking detectors Anderson (1933). Tracks with striking, surprising trajectories may exist in current experimental datasets, undiscovered only because they are invisible to current algorithms.
@@
Most traditional track-finding algorithms Collaboration (2017) perform alternating track parameter fitting and iterative hit finding, which requires knowledge of the explicit parametric form of the track, and so are fundamentally unsuited to model-agnostic tracking. However, recent developments in machine learning such as graph neural networks (GNNs) Bronstein et al. (2017) allow for track-finding algorithms where the form of the track is never specified explicitly, only implicitly via the sample used to train the algorithm Ju et al. (2020a, 2021a). While not the original goal of ML tracking, this new approach separates the finding and fitting steps Caillou et al. (2024, 2022a), and therefore allows for finding non-helical trajectories when trained on such examples Sha et al. (2024). Crucially, it may also open the door to model-agnostic tracking, if the broad space of possible trajectories can be described implicitly via a training sample, and the tracker learns to generalize beyond the specific training examples.
@@
In Section II, the simulation of tracks is described, including the detector geometry and model-agnostic track generation. In Section III, the tracking pipeline is described. Section III details the reconstruction performance on various sets of tracks, including background rejection and tests of generalization. Section IV discusses the results and next steps.
@@ II.1 Detector Geometry
Particle energy loss and hit location uncertainty are neglected, but similar work did not show strong sensitivity Sha et al. (2024).
@@ II.2 SM Background
Madgraph5 Alwall et al. (2014) is used to generate $pp\rightarrow t\bar{t}$ events, which are showered and hadronized using Pythia8 Bierlich et al. (2022). Initial velocity and momenta are used to propagate Standard Model particles in helical trajectories under the assumption that there is no energy loss or secondary scattering.
All events use a pileup of $\mu=1$. The impact of pile-up on performance is not yet known, but in other studies the same GNN pipeline performed well in high pile-up conditions similar to those expected at the high-luminosity LHC Ju et al. (2021b). Such performance may extend to the case of generalized smooth tracks, or it may be more susceptible to confusion in such hit dense environments. A rigorous study is out of scope for this initial proof-of-principle.
@@ II.3 Non-Helical Track Generation
The model-agnostic search aims to be sensitive to a very broad range of possible new particles with unexpected trajectories through the detector. However, it cannot accept every possible trajectory, or all sets of hits would be candidates. Instead, particle tracks are required to follow smooth paths, those which can be parameterized by a curve, $\vec{x}(t)$ where all orders of spatial derivatives with respect to time, $\frac{d^{n}x}{dt^{n}}$, exist and are finite, which accepts all particles with physical paths, where the net force and its derivatives, $m\frac{d^{n}\ddot{x}}{dt^{n}}$, are physical [^1] While this may be untrue for idealized or simplified systems which exhibit discontinuities in external forces, fully physical classical systems disregarding quantum effects such as decays, are subject to classical field interactions, which are infinitely differentiable with respect to position and time Landau and Lifschits (1975).. It rejects candidates whose paths are discontinuous or kinked. This method of track generation narrows our analysis by disregarding non-prompt decays, and particles that scatter off of detector material. While many particles exhibit non-smooth behavior in detector environments, a smoothness assumption provides a theoretically sound foundation for demonstrating this proof-of-concept analysis.
Without assumption about the underlying physical theory, three-component spatial curves can be fully described in a Fourier basis by three sequences of Fourier amplitudes, $(a_{n},b_{n},c_{n})$ defined in a Cartesian basis. Such paths, however, are not necessarily smooth unless each sequence approaches zero as $n\rightarrow\infty$ at least as fast as a Schwartz function f(n)f(n)Stein (2003); Hormander (1990); Reed (1980). Explicitly, smoothness is guaranteed if $a_{n},b_{n},c_{n}\leq f(n)$ for all $n$.
@@
For demonstration, non-helical tracks with a small number of frequencies are also prepared, by sampling from finitely many frequencies in Fourier space, where each mode has a maximum amplitude of 60cm. An $N$-frequency track is defined as the collection of points where a function $\vec{g}_{N}({t})=\sum_{i=1}^{3}\sum_{n=-N}^{N}c_{ni}\exp(\frac{2\pi int}{\lambda})\hat{x_{i}}$ intersects the detector layers, where $\hat{x_{i}}$ is a cartesian basis, and the $c_{ni}$ are all randomly generated with a uniform distribution in the interval $[0,60]$ cm, and $\lambda$ is set equal to the radius of the detector which is a choice further discussed in App. .3.
@@ III Tracking
Many tracking algorithms, such as a Kalman Filter, assume a parametric track path and ionization and interaction properties typical of known particles. An initial set of hits allows for initial estimates of the track parameters, which in turn allows for propagation of the track to subsequent detector surfaces and a calculation of a window within which additional hits should be considered. This hit-by-hit approach naturally weaves together track finding and fitting. Machine learning can be used in a variety of ways to perform track findingDuarte and Vlimant (2022); DeZoort et al. (2023); Thais et al. (2022). For example, ML can be used to propagate initial tracks in a Kalman Filter Heinrich et al. (2024), or more broadly to select the next hit Vaage (2022); Kortus et al. (2023).
The machine-learning-based approach used below does not require an explicit model of the trajectory. Instead, the relationships between hits on the same track are implicitly defined through the example tracks in the training sample. A track is represented as a directed graph, with hits as nodes. Track-finding pipelines, such as Exa.TrkX, which use graph neural networks achieve high track-finding efficiencies in a broad variety of particle experiments, even under high pile-up conditions Caillou et al. (2022b); Verma and Jena (2020); Hewes et al. (2021); Drielsma et al. (2021); Akram and Ju (2022); Jia et al. (2024); Biscarat et al. (2021); Zdybal et al. (2024); Andrews et al. (2021); DeZoort et al. (2021); Correia et al. (2024); Elabd et al. (2022). The Exa.TrkX pipeline first performs graph construction, where hits are linked into a graph, followed by edge classification, where graph edges are evaluated as their likelihood to connect two hits which belong to the same track, followed by graph segmentation, where edges are pruned and collected hit connections form track candidates. The Exa.TrkX pipeline employed here relaxes the requirement that hits are ordered in time according to their distance from the vertex, following Ref. Sha et al. (2024). Graph edges are evaluated by a metric learning model, which learns a mapping from the physical space of hits to an abstract latent space, such that hits on the same track are near each other and hits from other tracks are more distant.
The Exa.TrkX pipeline employed requires information about the hit ordering. Traditionally, the assumption is made that hit ordering follows the layers from inside out. As tracks may change direction, in this application the true hit ordering is used following Ref. Sha et al. (2024). For SM tracks, this has little impact as backscattering toward the production vertex occurs infrequently, and the fake rate remains low.
@@
Table 2: Shown is the track finding efficiency for non-helical tracks when training and testing are performed on disjoint spaces as described in Section III.2. Training and testing both have non-helical tracks and SM background
Train Test Efficiency Fake Rate
@@
III.1 Baseline tests
@@
In practice, non-helical tracks are expected to be found among many SM tracks. In the next stage, SM tracks are added to the training sample and given the same labels [^2] In previous work which attempted to learn one specific non-helical track pattern Sha et al. (2024), SM tracks were labeled as fake to suppress their reconstruction. In contrast, here the aim is to learn all possible smooth tracks. Efficiency is lower in the presence of many background tracks, as expected, but remains strong; see Tab. 1. Even in the presence of SM backgrounds, the tracking pipeline can efficiently find non-helical tracks.
III.2 Generalization tests
The training sample is necessarily finite, and is constructed using a finite number of Fourier modes from one of many Schwartz functions. An important question is whether the tracker has learned the general properties of smooth tracks, allowing it to generalize outside of its training sample. Previous work Sha et al. (2024) showed hints that the tracking pipeline may be capable of generalization, learning to reconstruct tracks which fall outside of the space described by the training set.
@@ III.2 Generalization tests
In one case, training on Schwartz space 5 and testing on 6-5, gives poor performance. The testing set has very long tracks, as shown in Figure. 5, which makes reconstruction under the double matching criterion very difficult. Generalization is possible, but not universal or guaranteed.
III.3 Dependence on Training Functions
@@
III.4 Results
The efficiency of the pipeline when trained using Schwartz function 19 is shown in Tab. 4 for a selection of testing samples. The training sample size has been increased to 65K events, each with one non-helical track and SM tracks from one SM event. This pipeline has a strong capacity to reconstruct many varieties of smooth tracks when correctly trained, while maintaining a low fake rate.
@@ III.4 Results
III.5 Quirk Reconstruction
The non-helical track finding pipeline is applied in this Section to a specific BSM theory, quirks Kang and Luty (2009) which leave non-helical tracks. A measure of the pipeline’s ability to recover physically motivated signals is provided by this Section.
Training is again performed on 65K events with a combination of non-helical tracks from Schwartz function 19 and SM tracks as in Section III.4, and testing is performed on 1K events containing both SM and quirk tracks. Following the quirk generation scheme of Sha et al. (2024), an overall efficiency of 32 percent on quirk tracks with 0 fakes is reported. The reconstruction efficiency versus the $p_{\mathrm{T}}$ of the quirk pair, the opening angle between pair produced quirks, the oscillation amplitude, and the oscillation period is displayed in Figure 8. Examples of reconstructed quirk tracks are shown in Figure  7.
Figure 7: Examples of reconstructed (predicted) quirk tracks
Figure 8: Shown is the reconstruction efficiency on quirks with respect to $p_{\mathrm{T}}$, oscillation period, amplitude, and opening angle
III.6 Rejection of Standard Model tracks
@@
Helical tracks can be rejected by fitting their points to a helical trajectory and evaluating the goodness of fit, via the $\chi^{2}$. For each track, an initial rapid estimation of the helical parameters is made Hansroul et al. (1988), followed by an exploration of parameters values in the vicinity, and refinement with the Nelder-Mead method Nelder and Mead (1965) implemented in scipy.minimize Virtanen et al. (2020). The distribution of the $\chi^{2}$ for SM and non-helical tracks is shown in Figure  9. While some SM tracks will have larger $\chi^{2}$, and some non-helical tracks have low $\chi^{2}$, there is a significant population of non-helical tracks with large $\chi^{2}$ and which are striking by eye. Several examples of fitting helices to hits left by non-helical tracks are shown in Figure . 10.
@@ V Conclusions
Acknowledgements
The authors are grateful to Makayla Vessella and Paata Ivanisvili for helpful discussion. LC and DW are supported by the DOE Office of Science.